williams2202:

The area of a rectangular plot 24 feet long and 16 feet wide will be doubled by adding an equal distance to each side of the plot. What is the distance added to each side?

jhonyy9:

the actual area A = 24×16 = ? doubled area 2A= (24 +x9)(16+x) = ? just you need solve this for x